a trucker traveling along a highway encounters road construction so that traffic is down to single lane. After driving 30 minutes, the road construction ends and the trucker is able to increase his speed by 64 mph for 4.5 hours. if his entire trip is 303 miles, what was his speed in the road construction?
Answer by jndarrell(58)

Create a D=RT box chart and fill in from below:
Instead of your typical car vs plane scenario, you are using "w/ traffic and w/o traffic"
With:
Rate = x mph
Time = 0.5hr
W/O:
Rate = 64mph
Time = 4.5hrs
In your scenario you are provided with a total distance. Whenever this is given, you will take the sum of each's R*T and set them equal to the total distance...
0.5x + 64(4.5) = 303
0.5x + 288 = 303
Subtract 288 from each side in the first step to get x alone
0.5x = 15
Divide each side by 0.5 to get x alone
x=30
Two common oops at the end
1. Remember to always report your units (30mph)
2. Double check what they are asking you to find.
Check your work by taking the found value of x and plug it into the d=RT equation we set:
0.5(30)+64(4.5)=303
Work it through!
:)